OpenStudy (anonymous):

Joshni threw a ball and its height is represented by the function h(t) = -16t2 + 30t + 5. Explain how she can determine the maximum height the ball reached.

OpenStudy (happyjack411):

s = -16t² + 30t + 5 Now, v = ds/dt => -32t + 30 = 0....at maximum height so, t = 30/32 => 15/16 seconds Then, s = -16(15/16)² + 30(15/16) + 5 => s = 19.0625 feet...the maximum height reached.